OpenStudy (anonymous):

Solve for Y(s)? Y(s)(s^2+s+5/4)=1/s-e^(-pi*s)/s+s.

OpenStudy (tkhunny):

It seems you have forgotten your Order of Operations. Can you please make another attempt and provide better information. The right hand side is anybody's guess. Hint: s^2 + s + 5/4 = \(s^{2} + s + \dfrac{5}{4}\) (s^2 + s + 5)/4 = \(\dfrac{s^{2} + s + 5}{4}\)